OpenStudy (anonymous):

a bus is travelling with 48 passengers. when it arrives at a stop, 'x' passengers get off & 3 get on. at the next stop half the passengers get off & 7 get on. there are now 22 passengers. find 'x'.

OpenStudy (unklerhaukus):

48 48 - x +3 (48 - x +3)/2 +7 (48 - x +3)/2 +7 = 22

OpenStudy (compassionate):

You beat me too it!

OpenStudy (anonymous):

then how do we do it?

OpenStudy (unklerhaukus):

solve the equation for x \[\frac{48 - x +3}2 +7 = 22\]

OpenStudy (unklerhaukus):

so do this , minus 7 form both sides, multiple both sides by 2 ... ....

OpenStudy (anonymous):

can we do\[\frac{ 48-x+3 }{ 2}=15\] ?

OpenStudy (anonymous):

yes you can do .. -> 48-x+3=30 ...

OpenStudy (unklerhaukus):

keep going...